OpenStudy (anonymous):

how do you find the equation of the line given:m=-2;f(1)=2

OpenStudy (amistre64):

y = mx -m(1) + 2

OpenStudy (amistre64):

the most common way is to start with a slope definition and work your way to a slope intercept

OpenStudy (amistre64):

but the point slope form is just as valid an equation for a line as any other

OpenStudy (amistre64):

to which: y-Py = m(x-Px) y - 2 = -2(x-1) in this case